# Egocentric Speaker Classification in Child-Adult Dyadic Interactions: From Sensing to Computational Modeling thanks:

### v1

@@ Abstract
Autism spectrum disorder (ASD) is a neurodevelopmental condition characterized by challenges in social communication, repetitive behavior, and sensory processing. One important research area in ASD is evaluating children’s behavioral changes over time during treatment. The standard protocol with this objective is BOSCC, which involves dyadic interactions between a child and clinicians performing a pre-defined set of activities. A fundamental aspect of understanding children’s behavior in these interactions is automatic speech understanding, particularly identifying who speaks and when. Conventional approaches in this area heavily rely on speech samples recorded from a spectator perspective, and there is limited research on egocentric speech modeling. In this study, we design an experiment to perform speech sampling in BOSCC interviews from an egocentric perspective using wearable sensors and explore pre-training Ego4D speech samples to enhance child-adult speaker classification in dyadic interactions. Our findings highlight the potential of egocentric speech collection and pre-training to improve speaker classification accuracy.
Index Terms:
@@
I Introduction
Autism spectrum disorder (ASD) [1, 2] is a neurocognitive developmental disorder that describes individuals with impairments in social communication and interactions along with repetitive and restrictive behaviors and interests. Even though ASD may be diagnosed at any age, many critical signs and symptoms of ASD can be observed early in life. Specifically, recent studies report that 1 in 36 children in the USA are diagnosed with ASD [3]. Children with ASD may face challenges in acquiring language skills and communicating with others, which can increase the risk of developing mental disorders such as anxiety and depression. Although there is limited knowledge about the causes of ASD, recent studies show that early treatment can lead to substantial improvement in a child’s development [4].
Evaluating the behavior changes over time during treatment involves clinicians conducting child speech and behavior analysis. Language samples are typically acquired through established clinical protocols involving semi-structured dyadic interactions between a child and clinicians to evaluate their communication abilities and behaviors. One prominent protocol of this kind is Brief Observation of Social Communication Change (BOSCC) [5] for tracking changes in social and communicative skills during treatment. These clinical observations aim to elicit spontaneous responses from children under different circumstances to obtain the effectiveness of treatment.
@@ I Introduction
  • We designed and conducted one of the earliest studies that collect egocentric speech signals using a wearable audio sensing device in child-adult interactions in a clinically relevant context.
@@
  • Instead of relying on existing pre-trained models for child-adult speech classification, the proposed work integrates a pre-training stage using the egocentric speech data collected from the Ego4D.
@@
  • We perform extensive fine-tuning experiments on the dataset collected from egocentric speech samples involving child-adult interactions following clinical observation protocol. Our experimental results demonstrate the effectiveness of pre-training for egocentric sensing in improving child-adult speech classification.
II Egocentric Audio Sensing
What sensing technology should we use? In this study, we propose to collect egocentric speech data from child-adult dyadic interactions using wearable sensors. Audio sensing technologies have evolved significantly in the last two decades, transitioning from cumbersome phonographs to hand-held microphones and portable lapel mics. Many traditional systems, however, remain bulky, expensive, and lack scalability. Recent technological advances in wearable sensors have enabled the emergence of applications for capturing social communications from an egocentric perspective in naturalistic settings [20, 21, 22]. For example, by using the miniature mics on these devices, a digital recorder [23, 24, 25, 21] can overcome some of the drawbacks of the traditional systems, capturing day-long audio data centered around a person. In this work, we prototype the experiments following a recently proposed wearable audio-sensing solution called TILES Audio Recorder (TAR)[21]. We would highlight that this wearable solution applies to both controlled and naturalistic settings.
How to design egocentric audio sensing experiments using TAR? The primary motivation for us to adopt TAR is its design and optimization for Android platforms, open-source nature, and compatibility with various form factors, including small, lightweight, and budget-friendly smartphones like the Jelly and Atom  [^1] https://www.unihertz.com/. We modify the original functionality in TAR to collect raw audio rather than low-level acoustic descriptors during child-adult dyadic interactions. Specifically, these child-adult interactions are conducted following the BOSCC protocol. In the experiment, both the child and the examiner wore TAR devices throughout the sessions. The child used the lightweight Jelly version for comfort, while the examiner wore the Atom, as shown in Figure 1. It is worth noting that we observed a better audio recording quality using Atom sensors than Jelly sensors.
Data Collection The data collection is a part of the extension to the study described in [26]. Children who participated in this study were instructed to wear the Jelly version of the TAR during their visit to conduct BOSCC interviews. The data in this work includes ten unique BOSCC sessions from 10 children between the ages of 2-7 years old. Particularly, among all children, six are male. Given that this study is an early effort to pilot egocentric audio sensing involving child-adult interactions, this cohort of data collection includes only three children with ASD.
III Egocentric Child-Adult Speaker Classification
III-A Pre-training The Backbone Model
Previous works have reported promising child-adult speaker classification results leveraging the recently developed speech foundation models. However, most existing speech foundation models are trained with speech recordings in a non-egocentric perspective. To improve the ability of these pre-trained speech models for egocentric speech understanding, we decided to perform a pre-training stage on a large-scale egocentric speech dataset called Ego4D.
Ego4D Dataset The Ego4D dataset is a popular dataset that has gained substantial interest in recent years. This multimodal dataset contains egocentric audio, videos, and IMUs from diverse geographic coverage, environments, and populations. The dataset includes over 3,500 hours of video data from 932 unique participants in 9 countries.
Voice Activity Detection As most audio data from the Ego4D dataset are unlabeled continuous naturalistic recordings, there exists a substantial amount of silence or environmental noises in these data. Therefore, we decide to preprocess the audio to obtain speech-only segments for the pre-training. Specifically, we leverage the voice activity detection model from pyannote [27] to filter out speech segments from long, continuous egocentric audio recordings.
@@ III-A Pre-training The Backbone Model
Figure 2: Pretraining framework of proposed modeling architecture using Ego4D speech data. The Ego4D audio samples were first processed by VAD.
@@
Wav2vec 2.0 Pre-training In this work, we primarily investigate the wav2vec 2.0 pre-training [28] on Ego4D speech samples. As shown in Figure 2, the wav2vec 2.0 model consists of convolutional encoders followed by transformers. The speech data $X$ is first mapped to latent feature $Z$, and the transformers model $Z$ to the contextualized representation $C$. Moreover, the latent feature $Z$ is discretized to $Q$ with a quantization module to form the contrastive objective. It is worth noting that the quantization module leverages the Gumbel Softmax operation to generate $Q$. The primary objective in wav2vec 2.0 pre-training is the contrastive object to distinguish the true quantized latent speech and its associated contextual representations after masking. More details regarding wav2vec 2.0 pre-training can be found in [28]. We want to mention that we initialized the model weights with the wav2vec 2.0-base model weights before further pre-training with speech samples filtered from the Ego4D dataset.
Figure 3: Fine-tuning architecture for child-adult speaker classification.
@@
III-B Speaker Classification Modeling
Once we complete the pre-training on Ego4D data, we perform the fine-tuning experiments following the fine-tuning model architecture from the previous work [10]. Specifically, we apply a linear layer to combine hidden outputs from all layers. The weighted output is then processed through a set of 1D convolutional layers with kernel sizes of 1, an average pooling layer, and classification layers of MLPs. We apply the ReLu activation functions within convolutional and MLP layers. The model architecture is demonstrated in Figure 3.
IV Experimental Details
IV-A Data Annotation
@@
IV-B Modeling Details
Ego4D Pre-training: The model weights of the pyannote VAD model were obtained from Huggingface for processing the Ego4D audio data. Since the child and adult dyadic interactions in our experiments were conducted using American English, we decided to filter out Ego4D audio data with the recording location within the United States. Moreover, those audios with a recording duration of less than $5$ minutes were also excluded from pre-training. After VAD, there were approximately 75,000 speech utterances for pre-training. We also initialize the pre-training model with the wav2vec 2.0 weights from Huggingface [^2] https://huggingface.co/facebook/wav2vec2-base. Our pre-training code was adopted from the github [^3] https://github.com/huggingface/transformers/tree/main/examples/pytorch/. Specifically, we choose a masking ratio of $50\%$, a learning rate of 0.0001, and a total training epoch of 40.
Speaker Classification: We apply a learning rate in $\{0.0002,0.0005\}$ for speaker classification training, with the CNN layers having an output channel size of 256. We perform the classification fine-tuning using two unique backbones: the wav2vec 2.0 base (W2V) and the Ego4D wav2vec 2.0 base (Ego4D W2V). We perform 5-fold cross-validation with two unique BOSSC sessions as a test fold and report the average performance across all folds. The reported metrics include accuracy, macro-F1, recall, and specificity. Here, adult and child labels are 0 and 1, respectively.
V Results
V-A Egocentric Child-Adult Speaker Classification
We begin by exploring whether pre-training the backbone model using egocentric speech samples improves child-adult speaker classification in an egocentric recording setup, as shown in Table I. In this experiment, the training data includes mono-channel speech samples from devices worn by both the child and the examiner. Compared with results reported in [10] (Macro F1 = 0.798), sampling audio with an egocentric setup helps to improve speaker classification. When tested on speech samples recorded from different sources (child or examiner), the findings suggest that pre-training the backbone model with egocentric speech samples consistently improves child-adult speaker classification. Specifically, we notice that the performance gains mainly originate from increased specificity, implying better adult speech classification. Given that most Ego4D speech samples come from adults, this highlights the benefit of pre-training on adult egocentric speech. It also points to the need for more egocentric speech samples from children to improve child speaker classification.
TABLE I: Child-adult speaker classification between two backbones: wav2vec 2.0 base (W2V) and a version further pre-trained with Ego4D speech samples (Ego4D W2V). The adult and child speaker labels are 0 and 1, respectively. The training data includes speech from both the child and the examiner.
Test Audio Backbone Acc F1 Recall Specificity
Child W2V $84.0$ $0.801$ $\mathbf{77.2}$ $87.8$
@@ V-A Egocentric Child-Adult Speaker Classification
V-B Impact of Training Recording Resource
We then compare the performances using a single source of recordings for training as shown in Table II. Specifically, we train the speech samples from either children or the examiner and test the models with the same or different recording sources.
Train/Test with the same recording source (Child/Child and Examiner/Examiner): Similar to the results presented earlier, we find that pre-training with the egocentric speech samples can consistently improve the child-adult speaker classification regardless of the source of training speech samples. Moreover, we note a higher child-adult classification performance when using the recording source from the examiner, confirming the better audio quality from the device worn by the examiner. We also observe that performance gains mainly come from improved adult speech classification.
Train/Test with different recording sources (Child/Examiner and Examiner/Child): Even when training and testing speech samples with varying sources of recording, we find that pre-training backbone models with egocentric speech samples still help child-adult speaker classification in these scenarios. However, we notice a consistent improvement in recall scores (child speech classification) when training and testing are used with speech samples from different recording sources instead of improved specificity.
TABLE II: Child-adult speaker classification results using different recording sources as training data. The adult and child speaker labels are 0 and 1, respectively.
@@ V-B Impact of Training Recording Resource
Figure 5: Performance comparisons at different training data ratios.
V-C Training Data Efficiency
Apart from leveraging complete annotated data for fine-tuning, we explore the efficiency of training data in child-adult speaker classification. Precisely, we compare two backbones in fine-tuning the child-adult speaker classification using $10\%,50\%,100\%$ of training samples. Results in Figure 5 demonstrate that Ego4D W2V consistently outperforms W2V across all training data ratios. These findings further support the effectiveness of integrating pre-training stages on egocentric speech samples to improve child-adult speaker classification sampled in the egocentric setup.
@@ V-D Generalizability to PEFT
V-E Dual Channel Modeling
Apart from classifying the child-adult speaker with the mono-channel input, we experimented with a concatenated input that combines speech samples from both devices. We would highlight that this modeling approach is similar to the dual-channel method proposed in the recent work [15]. Overall, the comparisons in Table III demonstrate that input with dual channels further improves child-adult speaker classification compared to the best-performing mono-channel approach. Compared with the previous work with conventional recording setup [10], dual-channel egocentric speech with Ego4D W2V backbone yields a 0.08 improvement in macro F1. Moreover, we can also observe a substantial performance increase in specificity using the Ego4D W2V backbone. This further implies the improved adult speaker classification that is associated with pre-training egocentric adult speech samples from the Ego4D dataset.
TABLE III: Child-adult speaker classification with dual channel modeling. The adult and child speaker labels are 0 and 1, respectively. The training samples include a two-channel speech from the child and the examiner.
@@ V-E Dual Channel Modeling
Dual W2V $87.9$ $0.855$ $84.7$ $91.1$
Ego4D W2V $\mathbf{89.7}$ $\mathbf{0.873}$ $\mathbf{79.6}$ $\mathbf{95.3}$
VI Conclusion and Future Works
In this work, we explore the use of wearable audio-sensing solutions to sample egocentric speech samples from child-adult interactions in clinical observations. We design the egocentric audio sensing experiments leveraging a previous wearable audio sensor successfully deployed in large-scale longitudinal studies in naturalistic settings. Moreover, we curate the data by annotating the speaker labels based on the prior work in [10]. On the other hand, we propose a modeling framework that begins with pre-training the backbone model with speech samples from the Ego4D dataset. Our extensive experiments show that speech samples collected in the egocentric setup can help pre-trained speech models to better classify child and adult speech. Moreover, pre-training the backbone model with egocentric speech samples can further improve the child-adult speaker classification.

### v2

@@ Abstract
Autism spectrum disorder (ASD) is a neurodevelopmental condition characterized by challenges in social communication, repetitive behavior, and sensory processing. One important research area in ASD is evaluating children’s behavioral changes over time during treatment. The standard protocol with this objective is BOSCC, which involves dyadic interactions between a child and clinicians performing a pre-defined set of activities. A fundamental aspect of understanding children’s behavior in these interactions is automatic speech understanding, particularly identifying who speaks and when. Conventional methods in this area heavily rely on audio recorded from a spectator perspective, and there is limited research on egocentric speech modeling. Here, we experiment with wearable sensors to perform speech sampling in BOSCC interviews from an egocentric perspective. Our findings highlight the potential of egocentric speech collection and pre-training to improve speaker classification accuracy.
keywords:
@@
1 Introduction
Autism spectrum disorder (ASD) [1, 2] is a neurocognitive developmental disorder that describes individuals with impairments in social communication and interactions along with repetitive and restrictive behaviors. Children with ASD may face challenges in acquiring language skills and communicating with others, which can increase the risk of developing mental disorders such as anxiety and depression. Recent studies show that early treatment of ASD can lead to substantial improvement in a child’s development [3]. Evaluating the behavior changes over time during treatment involves clinicians conducting child speech and behavior analysis. Language samples are typically acquired through established clinical protocols involving semi-structured dyadic interactions between a child and clinicians to evaluate their communicative abilities and behaviors. One prominent protocol is Brief Observation of Social Communication Change (BOSCC) [4] for tracking changes in social and communicative skills during treatment. These clinical observations aim to elicit spontaneous responses from children under different circumstances to obtain the effectiveness of treatment.
@@ 1 Introduction
  • We conducted one of the earliest studies that collect egocentric speech signals using a wearable audio sensing device in child-adult interactions in a clinically relevant context.
@@
  • Instead of relying on existing pre-trained models for child-adult speaker recognition, the proposed work integrates a pre-training stage using the egocentric speech from the Ego4D.
@@
  • We perform extensive fine-tuning experiments on the dataset collected from egocentric speech samples involving child-adult interactions following BOSCC protocol. Our results demonstrate the effectiveness of pre-training for egocentric sensing in improving child-adult speech classification.
@@
Figure 2: Our proposed pretraining framework. The Ego4D audio samples were first processed by VAD.
2 Egocentric Audio Sensing
Sensing technology In this study, we propose to collect egocentric speech from child-adult interactions using wearable sensors. Although audio sensing technologies have evolved significantly in the last two decades, many systems remain bulky, expensive, and lack scalability. Recent technological advances in wearable sensors have enabled the emergence of applications for capturing social communications from an egocentric perspective in naturalistic settings [19, 20, 21, 22]. For example, by using the miniature mics on these devices, a digital recorder [23, 24, 25, 20] can overcome many drawbacks of the traditional systems, capturing day-long audio data centered around a person. In this work, we prototype the experiments following a previously proposed wearable solution called TILES Audio Recorder (TAR) [20]. We would highlight that this wearable solution applies to both controlled and naturalistic settings.
Study Design The primary motivation for us to adopt TAR is its open-source nature and compatibility with various form factors, including lightweight and budget-friendly smartphones like the Jelly and Atom  [^1] https://www.unihertz.com/. We modify the original functionality in TAR to collect raw audio rather than low-level acoustic descriptors during child-adult dyadic interactions. Specifically, these child-adult interactions are conducted following the BOSCC protocol. In the experiment, both the child and the examiner wore TAR devices throughout the sessions. The child used the lightweight Jelly version for comfort, while the examiner wore the Atom, as shown in Figure 1. It is worth noting that we observed a better audio recording quality using Atom sensors than Jelly sensors.
Data Collection The data collection was a part of the extension to the study described in [26]. Children who participated in this study were instructed to wear the Jelly version of the TAR during their visit to conduct BOSCC interviews. The data includes ten unique BOSCC sessions from 10 children between the ages of 2-7 years old. Particularly, among all children, six are male. Given that this study is an early effort to pilot egocentric audio sensing involving child-adult interactions, this cohort of data collection includes only three children with ASD.
3 Egocentric Speech Modeling
3.1 Pre-training The Backbone Model
Previous works have reported promising child-adult speaker classification results leveraging the speech foundation models. However, most existing speech foundation models are trained with speech recordings in a non-egocentric perspective. To improve the ability of these pre-trained speech models for egocentric speaker prediction, we perform a pre-training stage on a large-scale egocentric speech dataset called Ego4D.
Ego4D Dataset The Ego4D dataset is a popular multimodal dataset that contains egocentric audio, videos, and IMUs from diverse geographic coverage, environments, and populations. The dataset includes over 3,500 hours of video data from 932 unique participants in 9 countries.
Voice Activity Detection As most audio data from the Ego4D are unlabeled continuous naturalistic recordings, there exists a substantial amount of silence or noise in these data. Therefore, we decide to preprocess the audio to obtain speech-only segments for the pre-training. Specifically, we leverage the voice activity detection model from pyannote [27] to filter out speech segments from continuous egocentric recordings.
@@ 3.1 Pre-training The Backbone Model
Wav2vec 2.0 Pre-training In this work, we primarily investigate the wav2vec 2.0 pre-training [28] on Ego4D speech samples. As shown in Figure 2, the wav2vec 2.0 model consists of convolutional encoders followed by transformers. The speech data $X$ is first mapped to latent feature $Z$, and the transformers model $Z$ to the contextualized representation $C$. Moreover, the latent feature $Z$ is discretized to $Q$ with a quantization module to form the contrastive objective. It is worth noting that the quantization module leverages the Gumbel Softmax operation to generate $Q$. The primary objective in wav2vec 2.0 pre-training is the contrastive object to distinguish the true quantized latent speech and its associated contextual representations after masking. More details regarding wav2vec 2.0 pre-training can be found in [28]. We initialized the weights with the wav2vec 2.0-base model weights before further pre-training with speech samples filtered from the Ego4D dataset.
Figure 3: Proposed fine-tuning architecture.
@@
3.2 Speaker Modeling
Once we complete the pre-training on Ego4D data, we perform the fine-tuning following the architecture from the previous work [9]. Specifically, we apply a linear layer to combine hidden outputs from all layers. The weighted output is then processed through a set of pointwise 1D convolutional layers, an average pooling layer, and MLPs for prediction. We apply the ReLu activation functions within convolutional and MLP layers. The model architecture is demonstrated in Figure 3.
4 Experimental Details
4.1 Data Annotation
@@
4.2 Modeling Details
Ego4D Pre-training: The model weights of the pyannote VAD model were obtained from Huggingface for processing the Ego4D. Since this study was conducted using American English, we decided to filter out Ego4D audio data with the recording location within the United States. Moreover, those audios with a recording duration of less than $5$ minutes were also excluded from pre-training. After VAD, there were approximately 75,000 speech utterances for pre-training. We initialize the pre-training model with the wav2vec 2.0 weights from Huggingface [^2] https://huggingface.co/facebook/wav2vec2-base. Specifically, we choose a masking ratio of $50\%$, a learning rate of 0.0001, and a total training epoch of 40.
Speaker Classification: We apply a learning rate in {2e-4, 5e-4} for speaker classification training, with the CNN layers having an output channel size of 256. We perform the classification fine-tuning using two unique backbones: the wav2vec 2.0 base (W2V) and the Ego4D wav2vec 2.0 base (Ego4D W2V). We perform 5-fold cross-validation with two unique BOSSC sessions as a test fold and report the average performance across all folds. The reported metrics include accuracy, macro-F1, recall, and specificity for the classification task. In the classification task, adult and child labels are 0 and 1, respectively.
5 Results
5.1 Egocentric Child-Adult Speaker Modeling
We begin by exploring whether pre-training the backbone model using egocentric speech samples improves child-adult speaker classification in an egocentric recording setup, as shown in Table 1. In this experiment, the training data includes mono-channel speech samples from devices worn by both the child and the examiner. Compared with results reported in [9] (Macro F1 = 0.798), sampling audio with an egocentric setup helps to improve speaker classification. When tested on speech samples recorded from different sources (child or examiner), the findings suggest that pre-training the backbone model with egocentric speech samples consistently improves child-adult speaker classification. Specifically, we notice that the performance gains mainly originate from increased specificity, implying better adult speech classification. Given that most Ego4D speech samples come from adults, this highlights the benefit of pre-training on adult egocentric speech. It also points to the need for more egocentric speech samples from children to improve child speaker classification.
Table 1: Child-adult speaker classification between two backbones: wav2vec 2.0 base (W2V) and a version further pre-trained with Ego4D speech samples (E4D W2V). The adult and child speaker labels are 0 and 1, respectively. The training data includes speech from both the child and the examiner.
Test Set Backbone Acc F1 Recall Specificity
Jelly-Child W2V $84.0$ $0.801$ $\mathbf{77.2}$ $87.8$
@@ 5.1 Egocentric Child-Adult Speaker Modeling
5.2 Impact of Training Recording Resource
We then compare the performances using a single source of recordings for training as shown in Table 2. Specifically, we explore the training with domain matching and domain mismatching. Specifically, domain matching indicates that the training and test have the same recording source, e.g., train and test both with recordings from the device the child is wearing (Child/Child). Instead, the domain mismatching represents the training and testing with recordings from different devices.
Domain Matching (Child/Child and Examiner/Examiner): Similar to the results presented earlier, we find that pre-training with the egocentric speech samples can consistently improve the child-adult speaker classification regardless of the source of training speech samples. Moreover, we note a higher child-adult classification performance when using the recording source from the examiner, confirming the better audio quality from the device worn by the examiner. We also observe that performance gains mainly come from improved adult speech classification.
Domain Mismatching (Child/Examiner and Examiner/Child): Even when training and testing speech samples with varying sources of recording, pre-training backbone models with egocentric speech samples still help child-adult speaker modeling in these scenarios. However, we notice a consistent improvement in recall scores (child speech classification) when training and testing are used with speech samples from different recording sources.
Table 2: Child-adult speaker classification results using different recording sources as training data. The adult and child speaker labels are 0 and 1, respectively.
@@ 5.2 Impact of Training Recording Resource
Figure 5: Performance at different training data ratios.
5.3 Training Data Efficiency
Apart from leveraging complete annotated data for fine-tuning, we explore the efficiency of training data in child-adult speaker modeling. Due to the limited space, we choose to only present speaker classification results. Precisely, we compare two backbones in fine-tuning the child-adult speaker classification using $10\%,50\%,100\%$ of training samples. Results in Figure 5 demonstrate that Ego4D W2V consistently outperforms W2V across all training data ratios. These findings further support the effectiveness of integrating pre-training stages on egocentric speech samples to improve child-adult speaker classification sampled in the egocentric setup.
@@ 5.4 Generalizability to PEFT
5.5 Dual Channel Modeling
Apart from modeling the child-adult speaker with the mono-channel input, we experimented with a concatenated input that combines speech samples from both devices. Similar to the PEFT experiment, we only explore the classification task. We would highlight that this modeling approach is similar to the dual-channel method proposed in [14]. Overall, the comparisons in Tab 3 demonstrate that input with dual channels further improves child-adult speaker classification compared to the best-performing mono-channel approach. Compared with the previous work with conventional recording setup [9], dual-channel egocentric speech with Ego4D W2V backbone yields a 0.08 improvement in macro F1. Moreover, we can also observe a substantial performance increase in specificity using the Ego4D W2V backbone. This further implies the improved adult speaker classification that is associated with pre-training egocentric adult speech samples from the Ego4D dataset.
Table 3: Child-adult speaker classification with dual channel modeling. The adult and child speaker labels are 0 and 1, respectively. The training samples include a two-channel speech from the child and the examiner.
@@ 5.5 Dual Channel Modeling
Dual W2V $87.9$ $0.855$ $\mathbf{84.7}$ $91.1$
Ego4D W2V $\mathbf{89.7}$ $\mathbf{0.873}$ $79.6$ $\mathbf{95.3}$
6 Conclusion and Future Works
In this work, we explore the use of wearable audio-sensing to sample egocentric speech samples from child-adult interactions in clinical observations. We design the egocentric audio sensing experiments leveraging a wearable audio sensor tested in large-scale longitudinal studies in naturalistic settings. Moreover, we curate the data by annotating the speaker labels based on our prior work in [9]. On the other hand, we propose a modeling framework that begins with pre-training the backbone model with speech samples from the Ego4D dataset. Our extensive experiments show that speech samples collected in the egocentric setup can help pre-trained speech models to better classify child and adult speech. Moreover, pre-training the backbone model with egocentric speech samples can further improve the child-adult speaker classification. Our future work includes comparing speaker diarization with previous child-adult speaker diarization models [31] and comparing with other speech foundation models.
7 Acknowledgements
We gratefully acknowledge support from the Simons Foundation (award number: SFI-AR-HUMAN-00004115-03, 655054). Some of our images are mixed from OpenMoji.